OpenStudy (anonymous):

Translate the phrase into a mathematical expression and then evaluate it. Thirty-six squared divided by six to the fourth power. The given expression in symbols is? Evaluated,the expression is equal to?

OpenStudy (anonymous):

\[\frac{ 36^2 }{ 6^4 }\]

OpenStudy (anonymous):

\[36^2 = 6*6*6*6 = 6^4\] So. \[\frac{ 36^2 }{ 6^4 } = 1\]

OpenStudy (anonymous):

1296/1296=1

OpenStudy (anonymous):

Yes!